Eric's Easter Eggs manufactures large chocolate eggs. The length of these eggs follows a normal distribution with mean = 13 inches and standard deviation = 1.8 inches.

If a sample of 40 eggs is selected, what is the probability that the sample mean length will be less than 12.5?

Must show ALL work
Answer #1

Solution :

Given that ,

mean = = 13

standard deviation = = 1.8

n = 40

=   = 13

= / n = 1.8 / 40 = 0.2846

P( < 12.5) = P(( - ) / < (12.5 - 13) /0.2846 )

= P(z < -1.76 )

Using the standard normal table

= 0.0392

probability = 0.0392